OpenStudy (anonymous):

83.3333 rounded to nearest tenth of a percent

OpenStudy (anonymous):

Would it be 83.3 or 83.33?

OpenStudy (anonymous):

The places after the point are tenths, hundredths, thousandths, etc.

OpenStudy (anonymous):

A tenth of a percent = 1/10 * 100 = 0.001 which is thousandths.

OpenStudy (anonymous):

83.3

OpenStudy (anonymous):

That's tenths.

OpenStudy (anonymous):

yea so its 83.3

OpenStudy (anonymous):

A tenth of a percent = 1/10 * 100 = 0.001 which is thousandths.